OpenStudy (anonymous):

voyager 2 traveled approximately 4.4 billion miles in 2.0 years. what was voyager 2's average speed, expressed in km/s?

OpenStudy (abmon98):

1 year = 730 days, right? In hours, that's 24hrs X 730 = 17520 hrs In minutes, that's 60 minutes X 17520 hrs = 105,120,0 minutes In seconds, that's 60 seconds X 105,120,0= 63,072,000 seconds Hundred - 100 Thousand - 1,000 Million - 1,000,000 Billion - 1,000,000,000 Trillion - 1,000,000,000,000 Quadrillion - 1,000,000,000,000,000 1 Mile = 1.690934 Km 4.4*10^9= x km x km=4.4*10^9*1.690934=7440109600 Average Speed=Total Distance/Total Time A.S=117.96 Km/second

OpenStudy (jfraser):

@Abmon98, I'm sure you mean in \(\large{2}\) years there are 730 days, not 1 year?

OpenStudy (abmon98):

oh i am sorry, thanks @JFraser :)